OpenStudy (anonymous):

@mathstudent55 Can you please help me?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

Thanl you..... What equation is solved by the graphed systems of equations? http://learn.flvs.net/webdav/assessment_images/educator_algebra2_v13/06_03_quiz/06_03_10.jpg

OpenStudy (anonymous):

7x + 3 = x + 3 7x - 3 = x - 3 7x + 3 = x - 3 7x - 3 = x + 3

OpenStudy (mathstudent55):

I can't see your figure. I get an error.

OpenStudy (mathstudent55):

Look at my figure below and compare it with your graph. Can you read point A? 

OpenStudy (anonymous):

It's down by a negative number

OpenStudy (mathstudent55):

Right. What is the negative number on the y-axis that that line goes through?

OpenStudy (mathstudent55):